OpenStudy (anonymous):

what is the period of the sinusoid given by y=-2sin((2pi/7)x)

OpenStudy (noelgreco):

\[y=a \sin(bx+c) +d\] This is the standard form of the sin function. \[T=\frac{ 2\pi }{ b } \] where T is the period of the function. "a" changes the amplitude only (vertical stretch). ""c" is the left/right phase shift. "d" moves the whole graph up or down.

OpenStudy (anonymous):

would it be 2pi/1?

OpenStudy (anonymous):

I would assume it would be 7

OpenStudy (danjs):

Following Noel post above. \[b=\frac{ 2\pi }{ 7 }\] Period is 2 pi divided by b

OpenStudy (anonymous):

after recalculating the problem

OpenStudy (danjs):

\[Period = \frac{ 2\pi }{ \frac{ 2\pi }{ 7 } } = \frac{ 2\pi }{ 1 }*\frac{ 7 }{ 2\pi } = \]

OpenStudy (danjs):

sin(2pi/7 x)

OpenStudy (anonymous):

thanks!

OpenStudy (noelgreco):

The coefficient of x in (2pi/7)x is 2 pi/7. Take a look at DanJS comment ten minutes ago, and you'll see that the period is 7 because the 2pis divide out